OpenStudy (anonymous):

I'm stuck!! I will reward help for medals. 5/12, 1/12, 1/3, 1/6 which one is the greatest? Please Please PLEASE! HELP!

OpenStudy (judygreeneyes):

Hello! To answer this question using fractions, you need to know about LCD's (least common denominators). Basically you have to convert all the fractions so that they have the same denominator. Of the denominators you have, 12, 12, 3 ,and 6, all of them can be multiplied to equal 12, so 12 can be your common denominator. You don't have to do anything to the first two fractions. To convert 1/3, you multiply the fraction by 4/4, so \[\frac{ 1 }{ 3 }x \frac{ 4 }{ 4 } = \frac{ 4 }{ 12 }\] The value of the fraction is the same, but now it is written in twelfths. Similarly, for 1/6, you need to multiply by 2/2 so you can get 12 in the denominator: \[\frac{ 1 }{ 6 }x \frac{ 2 }{ 2 }=\frac{ 2 }{ 12 }\] Now you can compare apples to apples! Which is larger? \[\frac{ 5 }{ 12}, \frac{ 1 }{12 }, \frac{ 4 }{ 12 }, or \frac{ 2 }{ 12} ?\]